OpenStudy (anonymous):

Which fraction is equivalent to eight over twenty? four over five two over ten two over five one over four

OpenStudy (anonymous):

2/5 I think

OpenStudy (nurali):

8/20 2/5

OpenStudy (anonymous):

Two over five. :3

OpenStudy (anonymous):

8/20 = 4/10 = 2/5

OpenStudy (anonymous):

two over five

hero (hero):

\[\frac{8}{20} = \frac{4 \times 2}{4 \times 5} = \frac{4}{4} \times \frac{2}{5} = \frac{2}{5}\]

OpenStudy (anonymous):

thanx everone not hero

OpenStudy (nurali):

Anytime.

OpenStudy (anonymous):

welcome